OpenStudy (anonymous):

find the quotient of 20x^2y-88xy/-4xy

OpenStudy (campbell_st):

just simply split the problem \[\frac{20x^2y}{-4xy} - \frac{88xy}{-4xy}\] now simplify for the quotient

OpenStudy (anonymous):

Yea ik its 5-22 but I don't know the rest

OpenStudy (campbell_st):

well look at the 1st fraction \[\frac{-4 \times -5 \times x \times x \times y}{-4 \times x \times y}\] what cancels, and what is left?

OpenStudy (anonymous):

-5 and x

OpenStudy (campbell_st):

yes and the 2nd fraction is \[\frac{-4 \times 22 \times x \times y}{-4 \times x \times y}\] what cancels..?

OpenStudy (campbell_st):

and what's left?

OpenStudy (anonymous):

22

OpenStudy (campbell_st):

thats correct... so when you combine the 2 fractions you get -5x + 22

OpenStudy (anonymous):

Ok